OpenStudy (anonymous):

Calculate the longitude of this parametric function, between 0 and Pi

OpenStudy (anonymous):

C = x(t) = \[2^{-t} * \cos t\] ; y(t) = \[2^{-t} * \sin t\] ; \[0 \le t \le \Pi \] Calculate longitude of C between 0 and Pi using the integral of \[\sqrt{d x^{2} + d y^{2}}\]

OpenStudy (anonymous):

Plase, help me. :S I can't solve the integral of \[\sqrt{d x^{2} + d y^{2}}\]

OpenStudy (amoodarya):

 can you write amd simplify that ?